Two factory plants are making TV panels. Yesterday, Plant A produced 6000 fewer panels than Plant B did. Five percent of the panels from Plant A and 4% of the panels from Plant B were defective. How many panels did Plant B produce, if the two plants together produced 690 defective panels?
Answer by ikleyn(52781)

.
Two factory plants are making TV panels. Yesterday, Plant A produced 6000 fewer panels than Plant B did.
Five percent of the panels from Plant A and 4% of the panels from Plant B were defective.
How many panels did Plant B produce, if the two plants together produced 690 defective panels?
~~~~~~~~~~~~~~~


x = the number of panels Plant B produced yesterday.


Plant A produced (x-6000) panels.


Total defective panels equation

    0.05*(x-6000) + 0.04x = 690


Simplify and find x

    0.05x - 300 + 0.04x = 690

         0.09x          = 690 + 300 = 990

             x                      = 990/0.09 = 11000.


ANSWER.  Yesterday, plant B produced 11000 panels.

Solved.